3 Due to erratic sales of its sole product-a high-capacity battery for laptop computers-PEM, Inc., has been experiencing financial difficulty for some time. The companys contribution format income statement for the most recent month is given below 4.5 points Sales (13,300 units x $30 per unit) Variable expenses Contribution margin Fixed expenses Net operating los:s 399,000 199,500 199,500 222,000 (22,500) eBook Required: 1. Compute the companys CM ratio and its break-even point in unit sales and dollar sales 2. The president believes that a $6,300 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $84,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the companys monthly net operating income? 3. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $37,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)? 4. Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by 0.50 cents per unit Assuming no other changes, how many units would have to be sold each month to attain a target profit of $4,500? 5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $51,000 each month. a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales. b. Assume that the company expects to sell 20,500 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.) C. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,500)? Print References
Answer #1

Requirement 1:

CM ratio=(contribution/sales)*100

=(199500/399000)*100= 50%

Break even point sales ($)=fixed cost /CM ratio

=222000/50%= $444000

Break even point in units= break even sales/seeking price per unit

= 444000/30=14800 units

Requirement 2:

Variable cost per unit= 199500/13300= $15 per unit

New sales= 399000+84000=483000

New fixed cost= 222000+6300= $228300

Sales=(fixed cost+ desired profit or loss)/CM ratio

483000= (228300+profit or loss)/50%

Profit or loss= $13200(profit)

Increase in income =13200+22500= $35700

Note: additional monthly advertising expenses are treated to be fixed expenses and hence are added to the existing fixed expenses. This assumption is maintained for the whole problem.

Requirement 3:

New seeking price per unit= 30*90%= $27 per unit

Number of units sold=13300*2=26600 units

Particulars amount ($)
A.A.Sal (26600*27) 718200
B.B.Variab expense (15*26600) 399000
C.Contribution (A-B) 319200
D.Fixed expenses 259000
Net operating income(C-D) 60200

Requirement 4:

New variable cost per unit=15+0.5= 15.5 per unit

Contribution margin per unit= 30-15.5= $14.5 per unit

Sales=( fixed cost+desired profit)/contribution per unit

=( 222000+ 4500)/14.5= 15621 units

Requirement 5 a):

New variable cost per unit=15-3= $12 per unit

New fixed expenses= 222000+51000= $273000

Contribution margin per unit=30-12= $18 per unit

Break even point= fixed cost/contribution per unit=273000/18

=15167 units ,ie , $455010 sales (15167*30)